OpenStudy (rainbow_rocks03):

A fruit stand had 36 oranges, 28 lemons, and 16 grapefruits. What was the ratio of oranges to the total pieces of fruit? A. 9 : 7 B. 1 : 2 C. 9 : 11 D. 9 : 20

OpenStudy (rainbow_rocks03):

SOMEONE PLZ PLZ PLZ HELP ME and tell me how you got that answer

OpenStudy (anonymous):

add together all the fruits 36+28+16

OpenStudy (anonymous):

oranges : total # of fruits

OpenStudy (anonymous):

then reduce

OpenStudy (rainbow_rocks03):

36+28+16=80

OpenStudy (anonymous):

yes and how many oranges do you have?

OpenStudy (rainbow_rocks03):

36

OpenStudy (anonymous):

okay so the ratio is 36:80. now simplify. now find the greatest common factor of those two numbers.

OpenStudy (rainbow_rocks03):

22?

OpenStudy (anonymous):

can 36 be divided by 22? nope

OpenStudy (rainbow_rocks03):

oh wait 4 right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so no divide 36 by 4 and 80 by 4.

OpenStudy (rainbow_rocks03):

36/4=9 and 80/4=20 so 9:20=D

OpenStudy (anonymous):

yeah that is correct :)